A 5.38 kg block is placed on top of a 9.24 kg bloc
Answer #1

Horizontal force =F=70.6N

Coefficient of kinetic friction =\mu =0.28

Acceleration due to gravity =g=9.8m/s^{2}

Total frictional force on 9.24 kg block =f_{s}

f_{s}=\left [ 0.28\times (5.38+9.24)\times 9.8+5.38\times 9.8\times 0.28 \right ]N=54.88N

Resultant force on 9.24kg block =F_{r}

F_{r}=F-f_{s}=(70.6-54.88)N=15.72N

Acceleration on that block =a

a=\frac{15.72}{9.24}m/s^{2}=1.7m/s^{2}

Hence a. is the right answer